Determine an equation in the form y=mx+b of a line satisfying the given condition:
throught the point (-5,-5) and parallel to the line 5y-2x=5

Answer by checkley71(8403)

5Y-2X=5
5Y=-2X+5
Y=-2X/5+5/5
Y=-2X/5+1 THIS LINE HAS A SLOPE OF -2/5.
NOW REPLACE THE X +Y TERMS IN THE LINE EQUATION [Y=mX+b] WITH -5,-5 & SOLVE FOR THE Y INTERCEPT (b).
-5=-(2/5)-5+b
-5=10/5+b
b=-5-2
b=-7 ANSWER FOR THE Y INTERCEPT.
THUS THE EQUATION IS:
Y=-2X/5-7